9 lbs. of mixed nuts containing 55% peanuts were mixed with 6 lbs. of another kind of mixed nuts that contains 40% peanuts. What percent of the new mixture is peanuts?
Answer by jorel1380(3719)

(9x.55)+(6x.40)=4.95+2.4=7.35
7.35/9+6=7.35/15=49% of the new mixture is peanuts..